OpenStudy (anonymous):

The times taken to assemble a clock at a factory are approximately normally distributed with a given mean U= 3 hr and standard deviation Q = 0.5 hr. What percentage of the times are between 2 hr and 4 hr? A)34% B)47.5% C)68% D)95%

OpenStudy (kirbykirby):

Let \(X\) be the time taken to assemble the clock. X is normally distributed, so you can standardize it using : \[Z = \frac{X-\mu}{\sigma} \] where \(\mu\) is the mean and \(\sigma\) is the standard deviation. SO the problem is basically asking for \(P(2 \le X \le 4)\). So you can do this: \[ P(2 \le X \le 4)=P\left(\frac{2-3}{0.5} \le \frac{X-3}{0.5}\le\frac{4-3}{0.5}\right)=P(-2 \le Z \le 2)\]

OpenStudy (kirbykirby):

Then you could - Use a standard normal distribution to find this probability. - Use a calculator/software to compute it (if you are not required to learn to use normal tables) - Use the empirical rule that says that x% of the data will be within 2 standard deviations of the mean.